# SeFENet: Robust Deep Homography Estimation via Semantic-Driven Feature Enhancement

### v1

@@ Abstract
1 Introduction
@@
In this paper, we propose a method named SeFENet, which enhances structural features by semantic features in semantic-driven module to achieve robust homography estimation under harsh environments. Different from structure information, semantic features capture high-level information to emphasize the details of objects and ignore the interference in the background caused by harsh conditions. Thus, we leverage the module designed for high-level perception tasks to guide the proposed network in extracting effective semantic features from degraded images, thereby achieving robust estimation. Furthermore, since the large gap between semantic and structural features, we introduce a Semantic-Guide Meta Constrains module optimized by a meta-learning strategy to mitigate the feature gap, ensuring they jointly enhance the model’s effectiveness. Meanwhile, considering the limitation of existing methods in capturing information from images in harsh environments, we introduce a hierarchical scale-aware mechanism to increase the receptive fields. So as to focus on more contextual information and understand the relationship of objects in the image. To summarize, our contributions are as follows:
@@
  • We introduce a hierarchical scale-aware feature extraction module. By increasing the receptive field of feature extraction, it enables the model to better comprehend the contextual information within the image scene while filtering out irrelevant background noise.
@@
2 Related Work
2.1 Homography Estimation
@@ 2.2 Meta-Learning
The concept of Meta-Learning is first proposed by [32]. Today, meta-learning has gradually been interpreted as giving machines the ability to learn to learn. Meta-learning has been applied in many fields [39, 10].  [36] proposes a meta-learning approach to create models that are less easy to drift, which is called catastrophic forgetting, when trained on different datasets constantly. A revised meta-learning framework proposed by [28] allows pre-trained networks to learn to adapt to new domains and enhance cross-domain performance. [44] introduces a new conditional diffusion-based meta-learning called MetaDiff. The optimization procedure of base learner weights from Gaussian initialization to target weights is modeled by MetaDiff in a denoising way.  [21] employs a meta-learning strategy during training to learn enhanced modality-agnostic representations even when only limited omnimodal samples are available. [37] proposes a new meta-learning regularization mechanism called Minimax-Meta Regularization, which uses inverse regularization in the inner loop and ordinary regularization in the outer loop during training. This continuously enhances the performance of the meta-learning algorithm.
3 Method
@@
The structure of SeFENet network is shown in Fig. 2, which contains 3 modules: Target Aware Homography Estimation Module (TAHEM), Semantic Extraction Module (SEM) and Semantic-Guide Meta Constrains (SMC). First, we input the source image $I_{\mathrm{s}}$ and target image $I_{\mathrm{t}}$ into pre-trained TAHEM to get homography matrix. The hierarchical scale perception module is embedded in TAHEM to expand the receptive field of feature extraction, so as to improve the robustness of the network in harsh environments. Then we input the ground truth $I_{\mathrm{gt}}$, the overlapping region of $I_{\mathrm{s}}$ and $I_{\mathrm{t}}$, into SEM to get semantic features. After that we feed this feature and the intermediate feature of TAHEM into SMC, implicitly enhance the ability of the proposed homography estimation network to extract semantic features through the alternately optimized meta-learning strategy, leading to realize the homography estimation of degraded environment. The specific details of key modules are as follows:
3.1 Hierarchical Scale-aware
@@
3.2 Semantic-Guide Meta Constrains
@@
3.3 Meta-Learning strategy
@@
$\theta_{\mathrm{SMC}}$ $=\theta_{\mathrm{SMC}}-\eta_{\mathrm{SMC}}\triangledown_{\theta_{\mathrm{SMC}}}\big{(}\mathcal{L}_{\mathrm{m}}(F_{\mathrm{se}}^{1/8},F_{t}^{1/8})$
@@
where $\theta_{\mathrm{SMC}}$ denotes the parameters before and after the SMC update. $\eta_{\mathrm{SMC}}$ is the learning rate of the SMC module. $\triangledown_{\theta_{\mathrm{SMC}}}\mathcal{L}_{\mathrm{m}}(F_{\mathrm{se}},F_{\mathrm{t}})$ can be calculated by:
$\triangledown_{\theta_{\mathrm{SMC}}}\mathcal{L}_{\mathrm{m}}(F_{\mathrm{se}},F_{\mathrm{t}})=\frac{\partial\mathcal{L}_{\mathrm{m}}(F_{\mathrm{se}},F_{\mathrm{t}})}{\partial\theta_{\mathrm{SMC}}},$ (7)
@@
$\theta_{\mathrm{TAHEM}}$ $=\theta_{\mathrm{TAHEM}}-\eta_{\mathrm{TAHEM}}\triangledown_{\theta_{\mathrm{TAHEM}}}\big{(}\mathcal{L}_{\mathrm{h}}(I_{s},I_{t},I_{gt})$
@@
$\triangledown_{\theta_{\mathrm{TAHEM}}}\big{(}\mathcal{L}_{\mathrm{h}}(I_{s},I_{t},I_{gt})+\lambda\sum\mathcal{L}_{\mathrm{m}}(F_{\mathrm{se}},F_{\mathrm{t}})\big{)}$
@@
4 Experiment
4.1 Implementation Details
@@
4.2 Dataset and Experiment Settings
@@
Figure 6: Visualization results of some traditional and deep learning methods on the VOC dataset, where the green polygon indicates the position of the source image in terms of the truth value on the target image. The red box indicates the prediction in the target image after different homography estimation algorithms. The closer the distance between the 2 polygonal boxes, the better the homography estimation. O+R is ORB+RANSAC, O+M is ORB+MAGSAC, L+R is LoFTR+RANSAC.
@@
Table 1: The point matching errors(PME), PSNR, SSIM, NCC of our method and other methods on large baseline CAHomo, where the best results are highlighted in bold and the second best results are underlined. For PME, a lower index value indicates better homography estimation. For PSNR, SSIM, and NCC, a higher index value signifies better quality of the generated image.
4.3 Evaluation on VOC dataset
@@
In Fig. 5, we compare the performance of 16 methods on the VOC dataset in normal conditions and 3 harsh environments. The results show that harsh environments make feature point extraction challenging, leading the convergence speed of the proportion of different ace images to 1 is much slower than the method based on deep learning. However, with the growth of ACE, our method can make the proportion of images with different ACE converging to 1 fastest.
@@
We visualized the homography estimation results in Fig. 6. Traditional methods such as ORB+RANSAC, ORB+MAGSAC, and LoFTR+RANSAC exhibit significant instability, failing to produce satisfactory results across all 4 scenarios. For other deep learning methods, the synthetic large baseline dataset also poses a challenge due to increased noise in harsh environments, resulting in a considerable deviation between their results and the ground truth. In contrast, our method shows a nearly perfect match between the red box and the green box, indicating highly accurate homography estimation.
4.4 Evaluation on Large-Baseline CAHomo dataset
@@
To demonstrate the robustness of our method in real environments, we test it using an improved benchmark dataset based on CAHomo from  [20]. We employ the same comparison method as the VOC dataset test. PME is used to evaluate the homography estimation results, while PSNR, SSIM, and NCC assess the quality of the generated images. The experimental results are presented in Table. 1, showing that traditional feature point extraction methods are often unstable, leading to awful outcomes. Except for the LoFTR-based point fetching method, PME for images generated by traditional methods is much lower in natural conditions compared to harsh environments. Most deep learning methods exhibit better robustness with images having large baselines. However, our method shows a clear advantage over others in both normal and harsh environments. Compared to the PME of the second-place MCNet, our model reduces PME by 55.5%, 54.8%, 41.7%, and 49.6% across the 4 cases respectively. Additionally, our method performs well in PSNR, SSIM, and NCC. Traditional image generation methods often produce subpar results due to the challenges associated with feature point extraction in harsh environments, resulting in equal values of PSNR, SSIM, and NCC for some methods under certain conditions. While our method’s SSIM only lags behind ORB+MAGSAC and ORB+RANSAC in the rain case, it ranks first in all other metrics across all conditions, with at least a 3% improvement over the second-place metrics.
@@
4.5 Ablation Study
Optimal number of iteration modules: Our model includes an iteration module that recurrently extracts and estimates features at each scale. To determine the optimal number of iteration modules, we experimented with $n=3,4,5,6,$ and $7$. Fig. 8(a) shows the MACE values after training in 3 different environments with varying numbers of iteration modules. Therefore, we set the number of iteration blocks to 5 in our model.
Figure 8: (a) represents the values of MACE for the homography estimation of the model in low-light, haze and rain environments when the number of iteration blocks ranges from 3 to 7. (b) shows the values of MACE for the model’s results under the 3 environments with and without fusion of semantic features or the hierarchical scale-aware.
Figure 9: The ablation of different iterations, where the green polygon indicates the position of the source image in terms of the truth value on the target image. The red box indicates the prediction in the target image.
Effectiveness of hierarchical scale-aware and Semantic extraction: To demonstrate the effectiveness of hierarchical scale-aware and semantic features, we conducted ablation studies by removing these 2 modules separately from the model and then training it without them. The results are shown in Fig. 8 (b). In the experiment testing the effectiveness of hierarchical scale-aware module, we use features extracted by the model without this module for correlation computation. As the result depicted in Fig. 8(b), the MACE value of the model goes worse noticeably when the hierarchical scale-aware module is removed.
To assess the impact of fusing semantic features, we excluded the SEM and SMC modules, using only the TAHEM for training. The experimental results, shown in Fig. 8(b), indicate that the network incorporating SEM and SMC significantly outperforms the network with only TAHEM across 3 harsh environments. This demonstrates semantic information plays a crucial role in homography estimation.

### v2

@@ Abstract
Index Terms:
Low-level computer vision, Homography estimation, Meta learning
I Introduction
@@
In this paper, considering the limitation of existing methods in capturing information from images in harsh environments, we introduce a hierarchical scale-aware mechanism to increase the receptive fields. So as to focus on more contextual information and understand the relationship of objects in the image. Meanwhile, with the multi-scale structure feature, we combine them with multi-scale semantic features and propose a method named SeFENet, which enhances structural features by semantic features, which make homography network concerns more detailed information in the homography estimation in semantic-driven module, so that the network can better deal with the edge information and achieve robust homography estimation under either normal environment or harsh environments. Different from structure information, semantic features capture high-level information to emphasize the details of objects and ignore the interference in the background caused by harsh conditions. Thus, we leverage the module designed for high-level perception tasks to guide the proposed network in extracting effective semantic features from degraded images, thereby achieving robust estimation. Furthermore, since the large gap between semantic and structural features, we introduce a Semantic-Guide Meta Constrains module optimized by a meta-learning strategy to mitigate the feature gap, ensuring they jointly enhance the model’s effectiveness. To summarize, our contributions are as follows:
@@
  • We propose a hierarchical scale-aware feature extraction module that enhances the model’s ability to capture multi-scale contextual information within an image scene. By expanding the receptive field during feature extraction, our approach effectively preserves essential details while filtering out irrelevant background noise, leading to more robust and accurate representations.
@@
II Related Work
II-A Homography Estimation
@@ II-B Meta-Learning
Today, meta-learning has been applied in many fields, [36, 37].  [38] proposes a meta-learning approach to create models that are less easy to drift, which is called catastrophic forgetting, when trained on different datasets constantly. A revised meta-learning framework proposed by [39] allows pre-trained networks to learn to adapt to new domains and enhance cross-domain performance. [40] introduces a new conditional diffusion-based meta-learning called MetaDiff. The optimization procedure of base learner weights from Gaussian initialization to target weights is modeled by MetaDiff in a denoising way.  [41] employs a meta-learning strategy during training to learn enhanced modality-agnostic representations even when only limited omnimodal samples are available. [42] proposes a new meta-learning regularization mechanism called Minimax-Meta Regularization, which uses inverse regularization in the inner loop and ordinary regularization in the outer loop during training. This continuously enhances the performance of the meta-learning algorithm.
In practical applications, meta-learning has increasingly been applied to semantic information research, with a primary focus on enhancing model robustness in complex and dynamic environments. [43] proposed a meta-learning-based text-driven visual navigation model using semantic segmentation and Word2Vec to extract visual and semantic features, enabling quick adaptation to new tasks through meta-reinforcement learning. [44] introduced Metabdry, a domain generalization method for named entity boundary detection, using meta-learning to simulate domain transfer and reduce performance differences. For few-shot semantic segmentation, [45] developed SML, which incorporates class-level semantic descriptions and ridge regression to compute prototypes, improving model performance with minimal data. Lastly, [46] applied few-shot learning and meta-learning techniques to point cloud data semantic segmentation, effectively addressing issues related to label scarcity and class imbalance. Furthermore, by combining meta-learning with open composite domain adaptation, [47] developed a new framework to address the open-label problem in semantic segmentation.
II-C Object Detection
Modern object detection models are typically divided into two categories. [48]. One set is region-based. The other is single-stage detectors, which do not require generating candidate regions. Instead, they transform the object detection task into a regression problem, significantly improving inference speed. Such as the YOLO framework [49] and SSD [50].
Currently, the single task of object detection is being increasingly integrated with other downstream tasks. In the context of low-light enhancement, [51] proposes an edge computing-driven, end-to-end framework for image enhancement and object detection under low-light conditions. This framework includes cloud-based enhancement and edge-based detection phases, significantly improving detection performance on edge devices while maintaining low latency in low-light environments. However, [52] introduces a plug-and-play image enhancement model, also based on an end-to-end framework, which highlights the model’s ease of use and high efficiency in low-light conditions. Additionally, for the video analysis, [53] proposes a lightweight object detection network in fast-paced sports environments that reduces computational resource consumption and enhances detection robustness through multi-scale feature fusion and other techniques. Within this domain, object detection is further explored in target tracking applications. [54] introduces a novel tracking-with-detection paradigm, where detection generates candidate bounding boxes in each video frame to improve the precision and robustness of target tracking. Recently, [55] introduced ByteTrack, which enhances multitarget tracking performance by extending the association of detection boxes, thereby preventing the loss of low-scoring targets. This approach further advances the field of multi-target tracking.
Object detection continues to see innovation in practical applications. [56] introduces a new harmony search algorithm based on computational intelligence for real-time target detection and tracking in video surveillance systems, improving detection accuracy and tracking stability while efficiently identifying multiple targets in video frames. In autonomous vehicles fields, [57] proposes a single-stage object detection framework based on YOLOv4 to improve detection accuracy. Additionally, it introduces an optimized network pruning algorithm to address the challenge of meeting real-time performance requirements given the limited computational resources of vehicle-mounted platforms, thus improving the speed of model inference. In the field of digital medicine, object detection is also extensively applied in medical image analysis to assist doctors in accurately locating and identifying disease boundaries. [58] combines YOLOv5 with deep neural networks to detect dislocations in vertebral columns. [59] uses the YOLO object detection algorithm to implement an automatic grading and diagnostic system for prostate cancer.
III Method
@@
The structure of SeFENet network is shown in Fig. 2, which contains 3 modules: Target Aware Homography Estimation Module (TAHEM), Semantic Extraction Module (SEM) and Semantic-guide Meta Constrains (SMC). First, we input the source image $I_{\mathrm{s}}$ and target image $I_{\mathrm{t}}$ into pre-trained TAHEM to get homography matrix. The hierarchical scale perception module is embedded in TAHEM to expand the receptive field of feature extraction, this approach enables the model to simultaneously capture both global and local information, allowing for a more comprehensive understanding of image content. Additionally, in challenging environments, it helps mitigate the impact of local noise, ensuring more robust and reliable results, so as to improve the robustness of the network in harsh environments. Then we input the ground truth $I_{\mathrm{gt}}$, the overlapping region of $I_{\mathrm{s}}$ and $I_{\mathrm{t}}$, into SEM to get semantic features. After that we feed this feature and the intermediate feature of TAHEM into SMC, implicitly enhance the ability of the proposed homography estimation network to The alternating optimization meta-learning strategy is employed to extract semantic features, enabling the model to capture multi-scale information while mitigating the impact of structural deformation in challenging environments. By leveraging semantic information to comprehend the contextual perspective relationships, the approach effectively reduces the likelihood of mismatches, thereby achieving robust homography estimation in degraded conditions. The specific details of key modules are as follows:
III-A Hierarchical Scale-aware
In this module, we adopt a hierarchical attention mechanism that allows the network to focus on features at different scales across multiple iteration rounds. In the large-scale attention mechanism, we emphasize the global features of the image, capturing its overall structure. On the other hand, the small-scale attention mechanism concentrates on finer-grained detail features, enabling the network to capture subtle image information. By learning both global and local features simultaneously, our network improves its ability to perform homography estimation, leading to enhanced performance.
@@ III-A Hierarchical Scale-aware
III-B Semantic-Guide Meta Constrains
In this section, we integrate semantic features into the homography estimation network. While the existing homography networks primarily focus on the structural relationships between images, incorporating semantic features enables the network to also account for the semantic information at the model’s edges. This enhancement helps the network reduce issues such as ambiguity, leading to more accurate and robust homography estimation.
@@ III-B Semantic-Guide Meta Constrains
Methods Normal Low-Light Haze Rain
PSNR SSIM NCC PSNR SSIM NCC PSNR SSIM NCC PSNR SSIM NCC
1) SIFT+RANSAC 11.604 0.421 0.538 23.758 0.657 0.341 11.933 0.653 0.181 11.507 0.398 0.487
2) SIFT+MAGSAC 11.341 0.417 0.523 23.739 0.658 0.341 10.934 0.645 0.124 10.789 0.425 0.458
3) ORB+RANSAC 11.719 0.464 0.478 23.756 0.659 0.342 11.453 0.653 0.148 10.937 0.559 0.433
4) ORB+MAGSAC 10.864 0.548 0.482 23.740 0.658 0.341 10.858 0.646 0.116 10.344 0.628 0.424
5) SPSG+RANSAC 11.357 0.400 0.523 25.531 0.629 0.478 13.215 0.651 0.489 11.366 0.336 0.516
6) SPSG+MAGSAC 11.326 0.404 0.529 25.280 0.652 0.504 12.694 0.654 0.465 11.152 0.329 0.512
7) LoFTR+RANSAC 11.110 0.419 0.501 24.713 0.641 0.464 12.203 0.638 0.386 10.979 0.335 0.485
8) LoFTR+MAGSAC 11.133 0.419 0.503 24.793 0.645 0.473 12.177 0.640 0.394 10.979 0.334 0.487
9) BasesHomo 19.028 0.557 0.750 31.519 0.856 0.752 34.760 0.887 0.831 19.240 0.395 0.697
10) HomoGAN 20.496 0.626 0.842 32.017 0.869 0.880 35.859 0.903 0.982 20.337 0.471 0.830
11) DHN 18.651 0.528 0.722 31.274 0.846 0.721 34.932 0.882 0.810 19.052 0.378 0.698
12) LocalTrans 17.815 0.480 0.735 31.192 0.840 0.807 34.742 0.881 0.971 18.671 0.368 0.739
13) RealSH 14.415 0.397 0.498 29.242 0.810 0.552 31.737 0.870 0.865 15.651 0.304 0.519
14) MCNet 17.564 0.577 0.830 31.602 0.864 0.916 34.295 0.898 0.992 18.774 0.518 0.871
15) Ours 20.445 0.623 0.892 32.998 0.887 0.935 36.612 0.912 0.991 20.939 0.508 0.896
TABLE I: The PSNR, SSIM, NCC of our method and other methods on large baseline VOC, where the best results are highlighted in red bold, and the second best results are highlighted in blue bold. For PSNR, SSIM, and NCC, a higher index value signifies better quality of the generated image.
III-C Meta-Learning strategy
To better integrate semantic and structural information, we employ a meta-learning strategy. Through the interplay between the inner and outer loops, this strategy enables the model to optimize both the homography network and the feature fusion network simultaneously. As a result, the feature fusion network can more effectively assist in homography estimation, leading to improved performance.
@@ III-C Meta-Learning strategy
$\theta_{\mathrm{SMC}}$ $=\theta_{\mathrm{SMC}}-\eta_{\mathrm{SMC}}\nabla_{\theta_{\mathrm{SMC}}}\big{(}\mathcal{L}_{\mathrm{m}}(F_{\mathrm{se}}^{1/8},F_{t}^{1/8})$
@@
where $\theta_{\mathrm{SMC}}$ denotes the parameters before and after the SMC update. $\eta_{\mathrm{SMC}}$ is the learning rate of the SMC module. $\nabla_{\theta_{\mathrm{SMC}}}\mathcal{L}_{\mathrm{m}}(F_{\mathrm{se}},F_{\mathrm{t}})$ can be calculated by:
$\nabla_{\theta_{\mathrm{SMC}}}\mathcal{L}_{\mathrm{m}}(F_{\mathrm{se}},F_{\mathrm{t}})=\frac{\partial\mathcal{L}_{\mathrm{m}}(F_{\mathrm{se}},F_{\mathrm{t}})}{\partial\theta_{\mathrm{SMC}}},$ (8)
@@
$\theta_{\mathrm{TAHEM}}$ $=\theta_{\mathrm{TAHEM}}-\eta_{\mathrm{TAHEM}}\nabla_{\theta_{\mathrm{TAHEM}}}\big{(}\mathcal{L}_{\mathrm{h}}(I_{s},I_{t},I_{gt})$
@@
$\nabla_{\theta_{\mathrm{TAHEM}}}\big{(}\mathcal{L}_{\mathrm{h}}(I_{s},I_{t},I_{gt})+\lambda\sum\mathcal{L}_{\mathrm{m}}(F_{\mathrm{se}},F_{\mathrm{t}})\big{)}$
@@
IV Experiment
IV-A Implementation Details
@@
IV-B Dataset and Experiment Settings
@@
Figure 6: The first row in each environment displays visualization results of some traditional and deep learning methods on the VOC dataset, where the green polygon indicates the position of the source image in terms of the truth value on the target image. The red box indicates the prediction in the target image after different homography estimation algorithms. The closer the distance between the 2 polygonal boxes, the better the homography estimation. The second row in each environment shows the error image, which is the difference between the warped image and the target image. The closer the error image is to the ground truth, indicating more accurate homography estimation. Here, O+R refers to ORB+RANSAC, O+M to ORB+MAGSAC, and L+R to LoFTR+RANSAC.
@@
TABLE II: The point matching errors(PME), PSNR, SSIM, NCC of our method and other methods on large baseline CAHomo, where the best results are highlighted in red bold, and the second best results are highlighted in blue bold. For PME, a lower index value indicates better homography estimation. For PSNR, SSIM, and NCC, a higher index value signifies better quality of the generated image.
IV-C Evaluation on VOC dataset
We conducted a comprehensive evaluation on the improved VOC benchmark dataset in normal conditions and 3 harsh environments. To assess the quality of the generated images, we utilized PSNR, SSIM, and NCC, as presented in Table. I. As shown in the table, our approach exhibits clear advantages. It consistently ranked among the top two in nearly all test environments and achieved the best performance across most metrics. In the Normal environment, although PSNR and SSIM are slightly lower than those of HomoGAN, the average difference is only 0.365%, making the gap negligible. In contrast, the NCC is significantly higher, surpassing HomoGAN by nearly 6%, demonstrating stronger feature consistency. In complex environments, with the exception of the NCC in the Haze environment and the SSIM in the Rain environment, most other metrics outperform all competing methods, achieving an average improvement of nearly 3% compared to the second-best approach. Notably, in Low-Light environments, our method exhibited absolute superiority, outperforming all other methods across all evaluation metrics, further underscoring its exceptional capability in feature point extraction under harsh conditions. Overall, our approach excels in preserving structural information and visual quality, demonstrating greater adaptability and robustness in both standard and challenging environments.
@@ IV-C Evaluation on VOC dataset
Then, in Fig. 5, we compare the performance of 16 methods on the VOC dataset in 4 environments. The results show that harsh environments make feature point extraction challenging, leading the convergence speed of the proportion of different ace images to 1 is much slower than the method based on deep learning. However, with the growth of ACE, our method can make the proportion of images with different ACE converging to 1 fastest.
@@
We visualized the homography estimation results in the first row of each environment in Fig. 6. Traditional methods such as ORB+RANSAC, ORB+MAGSAC, and LoFTR+RANSAC exhibit significant instability, failing to produce satisfactory results across all 4 scenarios. For other deep learning methods, the synthetic large baseline dataset also poses a challenge due to increased noise in harsh environments, resulting in a considerable deviation between their results and the ground truth. In contrast, our method shows a nearly perfect match between the red box and the green box,
Figure 8: The ablation of different iterations, where the green polygon indicates the position of the source image in terms of the truth value on the target image. The red box indicates the prediction in the target image.Iter = 5 is our method.
Meanwhile, we visualized the error between the warped image and the target image, as shown in the second row in Fig. 6. Under the normal environment, although the overall error is small, some methods, such as O+M, DHN, and HomoGAN, exhibit large errors when compared to the error map generated by the true values, especially in the areas around the target box. The predictions of these methods deviate from the position of the source image, resulting in misalignment errors. In contrast, the error map generated by our method is almost perfectly consistent with the true value. In harsh environments, due to feature occlusion and blurring caused by environmental noise, most methods show large errors. Notably, in rainy conditions, RealSH and MCNet exhibit the most significant errors, with serious image misalignment that almost completely deviates from the true values. However, in these complex environments, our method performs particularly well. Even when affected by noise, the resulting error maps are smaller and more consistent with the true value compared to those generated by other methods. In summary, compared to other methods, our model get better homography matrix various environments. There is no obvious misalignment in the error maps, and they remain highly consistent with the error maps generated by the true values, further confirming the accuracy of our method in homography estimation.
Figure 9: The features extracted from models with and without meta-learning strategies are visualized using heap maps in normal, low-light, haze and rain environments. Cool color areas indicate areas where features are of higher importance. Warm color areas represent areas with less feature information. ML means meta-learning.
IV-D Evaluation on Large-Baseline CAHomo dataset
@@
To demonstrate the robustness of our method in real environments, we test it using an improved benchmark dataset based on CAHomo from  [20]. We employ the same comparison method as the VOC dataset test. PME is used to evaluate the homography estimation results, while PSNR, SSIM, and NCC assess the quality of the generated images. The experimental results are presented in Table. II, showing that traditional feature point extraction methods are often unstable, leading to awful outcomes. Except for the LoFTR-based point fetching method, PME for images generated by traditional methods is much lower in natural conditions compared to harsh environments. Most deep learning methods exhibit better robustness with images having large baselines. However, our method shows a clear advantage over others in both normal and harsh environments. Compared to the PME of the second-place MCNet, our model reduces PME by 55.5%, 54.8%, 41.7%, and 49.6% across the 4 cases respectively. Additionally, our method performs well in PSNR, SSIM, and NCC. Traditional image generation methods often produce subpar results due to the challenges associated with feature point extraction in harsh environments, resulting in equal values of PSNR, SSIM, and NCC for some methods under certain conditions. While our method’s SSIM only lags behind ORB+MAGSAC and ORB+RANSAC in the rain case, it ranks first in all other metrics across all conditions, with at least a 3% improvement over the second-place metrics.
indicating highly accurate homography estimation.
Figure 11: Metrics—including MACE, PSNR, SSIM, and NCC—were computed from both the VOC dataset and the CAHomo dataset and evaluated across iterative blocks (3 to 7) under normal, low-light, haze, and rain conditions. Lower MACE and PME values indicate better performance, while higher PSNR, SSIM, and NCC values reflect improvement.
Figure 12: The metrics calculated from the VOC dataset and CAHomo dataset remain consistent with previous evaluations but are used here to compare the performance of our model with and without the fusion of semantic features or hierarchical scale-awareness. Lower MACE and PME values indicate better performance, while higher PSNR, SSIM, and NCC values reflect improvement.
@@
IV-E Ablation Study
Optimal number of iteration modules: Our model includes an iteration module that recurrently extracts and estimates features at each scale. To determine the optimal number of iteration modules, we experimented with $n=3,4,5,6,$ and $7$. The similar results in Fig. 11 illustrate how the performance evolves as the number of iterations changes, providing a clearer understanding of the model’s behavior under different configurations. The first row of Fig. 11 represents the VOC dataset, and the second row represents the CAHomo dataset. From Fig. 11 (a), it is clear that when the number of iterations is set to 5, both the MACE values and PME values reach their lowest points, indicating optimal performance. When the number of iterations exceeds 5, the MACE values and PME values increase, resulting in performance deterioration. For the PSNR, SSIM, and NCC indicators across the two datasets, although the trend is not as pronounced as the above two indicators, in most cases, these values all remain at a high level when iteration is 5, which is better than when iteration is increased. This further confirms the superior performance of the five-iteration model. From the perspective of the model network architecture, this suggests that as the number of iterations increases, the network can begin to focus more on irrelevant or noisy details, rather than the meaningful features required for accurate homography estimation. Furthermore, from a model generalization perspective, another key factor to consider is the computational cost associated with increasing the number of iterations. As iterations increase, training time grows, which demands more computational resources, but also raises the possibility of inefficiency. In summary, considering the balance between model performance and training cost, we set the number of iterations in the model to 5.